# Constrained Black-Box Attacks Against Multi-Agent Reinforcement Learning

### v1

@@ IV-C The Align attack
Our problem in Equation (7) is similar to the one in Equation (2), suggesting that we can use PGD. However, there is a subtle detail that must be considered: in Equation (2) we add the perturbation to the input, whereas in Equation (7) we add it to both the input and the output.
In light of these factors, our attack can be split into the following two phases, as summarized in Algorithm 1:
Phase One: data collection. The purpose of this step is to collect a dataset $\mathcal{D}$ to train the neural network $f_{\theta}$. This dataset is gathered by storing the observations of infected agents during test time. Once we have accumulated observations for a sufficient period $\mathcal{T}^{c}$, we train the neural network $f_{\theta}$. This network is trained only once, and its parameters are frozen after this phase. No attack is executed during this phase.
Phase Two: adversarial attack. In this phase, we launch our adversarial attack. At each time step $t$, we start by intercepting the current observations $\{o_{i,t}\}_{1,\dots,n}$. Next, we generate the corresponding perturbations using PGD. Our exact implementation of PGD is outlined in Algorithm 2. Following this, we inject these perturbations into the respective agents.
@@
IV-D A lighter Align attack
It is not always practical to compute and inject adversarial perturbations for each agent. In addition to avoiding detection, one of the main reasons is the associated computational cost. A straightforward approach for handling this is to limit the number of agents being attacked. Instead of targeting all the agents, we only attack a subset. This subset is not randomly chosen; rather, it is based on the trained network $f_{\theta}$’s loss. We select a subset $\mathcal{M}\subset\mathcal{N}$ of the agents, where $m=|\mathcal{M}|<n$ with the lowest loss J, as this indicates that most of the agents have aligned perceptions regarding the informations contained in the $m$ observations.
Formally the subset $\mathcal{M}$ is defined as follows:
$\mathcal{M}=\operatorname{argmin}_{\begin{subarray}{c}\mathcal{S}\subset\mathcal{N}\\ |\mathcal{S}|=m\end{subarray}}\sum_{i\in\mathcal{S}}\text{J}_{i}(\mathbf{o};\theta)$ (8)
where $\text{J}_{i}$ is the per-agent loss defined as:
$\text{J}_{i}(\mathbf{o};\theta)=\left(f_{\theta}(o_{-i})-o_{i}\right)^{2}$ (9)
IV-E Free misalignment attacks
Our goal is to leverage the alignment intuition without access to any internal information about the agents (e.g., observations, actions, policies, etc.). To illustrate our method, imagine a 2D game in which two agents must coordinate to catch an object in order to receive a reward. Each agent receives the $(x,y)$ coordinates of the goal object as its observation. To introduce misalignment into the agents’ observations, an adversary can attempt to push the agents in directions that are not aligned (not similar and different). In our simple example, an effective adversarial strategy is to push each agent in perpendicular directions (e.g., $\theta=90^{\circ}$ in Equation (10), similar to cosine similarity in natural language processing).
$\delta_{1}^{\top}\delta_{2}=\lvert\delta_{1}\rvert\lvert\delta_{2}\rvert\cos\theta=0$ (10)
This idea generalizes to high-dimensional observation spaces using vector orthogonality. Suppose we have $n$ agents, each receiving an observation of dimension $d$. To design a misalignment attack, we will generate a matrix $\boldsymbol{\delta}$:
$\boldsymbol{\delta}=\begin{bmatrix}\delta_{11}&\delta_{12}&\cdots&\delta_{1d}\\ \delta_{21}&\delta_{22}&\cdots&\delta_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ \delta_{n1}&\delta_{n2}&\cdots&\delta_{nd}\end{bmatrix}\in\mathbb{R}^{n\times d}$ (11)
That must satisfy two conditions:
  1. 1.
    Condition 1: The rows are orthogonal :
  2. 2.
    Condition 2: Satisfy the attack budget constraint
    $\|\delta_{i}\|_{\infty}\leq\epsilon\quad\forall i\in\mathcal{N}$ (13)
Generating such perturbations is non-trivial. A simple greedy approach that constructs $\boldsymbol{\delta}_{i}$ one at a time can easily satisfy the budget constraint (13), but it becomes difficult to guarantee orthogonality (12), especially in high-dimensional spaces. Another strategy is to sample a random matrix (e.g., from a normal distribution) and orthogonalize it using QR decomposition. While this guarantees orthogonality, it does not ensure the budget constraint is satisfied.
To generate perturbations that satisfy both constraints, we use partial Hadamard matrices. A full Hadamard matrix $\mathbf{H}\in\mathbb{R}^{d\times d}$ is a square matrix with entries in $\{+1,-1\}$ satisfying:
$\mathbf{H}\mathbf{H}^{\top}=d\,\mathbf{I}.$
By construction, its rows are orthogonal. A partial Hadamard matrix is obtained by selecting a subset of rows from a full Hadamard matrix. The resulting matrix $\tilde{\mathbf{H}}\in\mathbb{R}^{n\times d}$, with $n<d$, retains exact orthogonality between its rows.
It is clear that the following matrix satisfies both conditions (1) and (2):
$\boldsymbol{\delta}=\epsilon\times\tilde{\mathbf{H}}$ (14)
In our implementation, we generate Hadamard matrices using Sylvester’s construction. It is important to note that Hadamard matrices exist only for specific dimensions: when $d$ is a multiple of 4.
To bypass the requirement that $d$ must be a multiple of 4, we generate a full Hadamard matrix of size $\tilde{d}$, where $\tilde{d}$ is the largest power of two such that $\tilde{d}\leq d$:
@@ IV-E Free misalignment attacks
Figure 3: Illustration: Padded partial Hadamard matrix for $n=3,d=10$.
@@
$\lambda\left\|\boldsymbol{\delta}^{\top}\boldsymbol{\delta}-\boldsymbol{I}\right\|_{2}^{2}.$
We implemented these two approaches, but did not notice any substantial improvements. However, using the lightweight version of the Align attack from the previous section to select a subset of agents and injecting Hadamard perturbations proved very effective in our experiments.
V Experiments
TABLE II: QMIX Hyperparameters Used for SMAC and LBF benchmark
@@ V Experiments
TABLE III: MAPPO Hyperparameters for SMAC benchmark.
@@
TABLE IV: MAPPO Hyperparameters for LBF benchmark.
@@
TABLE V: MAPPO Hyperparameters for RWARE.
@@
tiny-4ag 0.0005 False False 10 25M
tiny-4ag-sr2 0.0008 False False 10 25M
tiny-4ag-sr3 0.0008 False False 10 25M
small-4ag 0.0005 False False 10 25M
small-4ag-sr2 0.0008 False True 10 25M
small-4ag-sr3 0.0008 False False 20 40M
The goal of this section is to validate the effectiveness of our attack and identify the factors contributing to its outcome. To this end, we address the following key questions:
Question 1: Are the attacks effective? We seek to confirm that the execution of our attacks successfully compromises the performance of collaborative MARL.
Question 2: How do our attacks compare to standard baselines? We want to compare to standard adversarial attacks and see if our attacks outperform existing ones.
Question 3: Do the attacks generalize across different MARL algorithms? We examine whether the attacks affect both value-based and policy-based algorithms.
Question 4: What factors influence the Align attack’s outcome? We investigate several factors that could impact the Align attack’s success, primarily focusing on the size of the collected data, the architecture of the neural network $f_{\theta}$, the number of attacked agents, and the number of iterations $K$ of the PGD algorithm.
In the next section, we outline the experimental tools that help us to answer the aforementioned questions.
V-A Experimental setup
In order to answer our questions, we carefully choose multiple environments from three different MARL benchmarks:
  • Level-Based Foraging (LBF) [20]: LBF is a grid-world environment where agents collect randomly scattered food. We select three tasks: fully observable, highly cooperative, and partially observable. In fully observable tasks, agents receive the attributes of all entities in the grid (position and level). In the highly cooperative setting ( with the “-c” flag), food can only be collected if all agents act on it simultaneously. These setups provide overlapping observations, making alignment easier to measure. Partially observable tasks are expected to be more challenging for our attacks. In these tasks, agents only see a limited area around themselves. We test two settings: a 2-square radius (“-2s”) and a 1-square radius (“-1s”). We use 10 LBF environments in total. Details are in Table VI.
  • Multi-Robot Warehouse (RWARE) [20]: RWARE is a partially observable multi-agent environment where robots are tasked with collecting requested shelves inside a warehouse. We test with three different levels of partial observability, where agents observe a $7\times 7$, $5\times 5$, or $3\times 3$ grid. This benchmark also allows us to experiment with observations of higher dimensionality, in contrast to LBF’s observations. See table VII for details.
  • StarCraft Multi-Agent Challenge (SMAC) [3]: SMAC is the most widely used benchmark in c-MARL. It provides partially observable environments and includes scenarios with many agents, which helps us evaluate how our algorithms scale with a large number of agents. Since SMAC tasks are combat scenarios, they allow us to test our algorithms in situations where agents may become unavailable (e.g., due to death in SMAC). This mirrors many realistic settings where we might lose access to the target agents. See table VIII for details
TABLE VI: Selected LBF environments: we select (FO) fully observable, (PO) partial observable, and (HC) highly cooperative tasks.
Environment FO PO HC $\text{dim}(o_{i})$ N
8x8-3p-2f 15 3
8x8-3p-2f-coop 15 3
1s-8x8-3p-2f 15 3
2s-8x8-3p-2f 15 3
2s-8x8-3p-2f-coop 15 3
10x10-4p-2f 18 4
10x10-4p-2f-coop 18 4
1s-10x10-4p-2f 18 4
2s-10x10-4p-2f 18 4
2s-10x10-4p-2f-coop 18 4
TABLE VII: Selected RWARE environments: we select sensory range (SR) of $3\times 3$, $5\times 5$ and $7\times 7$.
Environment PO SR $\text{dim}(o_{i})$ N
tiny-4ag $3\times 3$ 71 4
tiny-4ag-sr2 $5\times 5$ 183 4
tiny-4ag-sr3 $7\times 7$ 351 4
small-4ag $3\times 3$ 71 4
small-4ag-sr2 $5\times 5$ 183 4
small-4ag-sr3 $7\times 7$ 351 4
TABLE VIII: Selected SMAC environments: all environments are partially observable (PO)
Environment PO $\text{dim}(o_{i})$ N
27m_vs_30m 285 27
10m_vs_11m 105 10
MMM2 176 10
protoss_5_vs_5 92 5
terran_5_vs_5 82 5
zerg_5_vs_5 82 5
As our paper considers test-time attacks, we need to first have well-trained agents. We train the agents in each environment with QMIX  [26] and MAPPO [29] using the hyperparameters specified in tables II, III, IV , and LABEL:tab:mappo_rware_hyper. We use the Epymarl library [20] for training. We only report hyperparameters that are different from the default values found in Epymarl configurations. For the RWARE benchmark, we only train MAPPO agents, as QMIX performs poorly at this task [20]. Once the training is complete, we freeze the policy weights.
For each environment, the adversary starts by collecting a dataset of observations from 5,000 environment transitions (i.e., $\mathcal{T}^{c}=5000$). This dataset is subsequently used to train the neural network $f_{\theta}$. We test with three different architectures:
Adversarial baselines: We use the following baselines:

### v2

@@ 4.3 The Align attack
The main benefits of our formulation in Equation (6) are threefold. First, it is black-box and does not require access to policy parameters. Second, unlike prior attacks, it does not rely on any policy or value function to generate perturbations. Third, it enables the use of PGD attacks similar to those in Equation (1). However, a subtle but important distinction must be noted. In Equation (1), the perturbation is added only to the input, whereas in Equation (6), it is added to both the input and the output.
To summarize, our attack consists of two phases. In the first phase, we collect the observations of deployed agents for a sufficient period $\mathcal{T}^{c}$, which are then used to train $f_{\theta}$. No attack is executed during this phase. In the second phase, we intercept the current observations $\{o_{i}\}_{1,\dots,n}$, generate the corresponding perturbations, and inject them into the agents. A full and detailed pseudocode can be found in the appendix.
4.4 Targeted Align attack
Attacking all agents may be computationally expensive and may reduce the stealthiness of the attack. Therefore, we may be forced to only attack a subset of agents $\mathcal{M}\subset\mathcal{N}$ with $m=|\mathcal{M}|<n$. For a targeted attack, we choose the subset of agents whose observations are most mutually aligned; we interpret them as agents who are most likely to coordinate among each other, and we aim to undermine their collaboration. Formally, for a fixed number of targeted agents $m$, we seek the subset $\mathcal{M}$ that satisfies the following:
$\mathcal{M}=\operatorname*{arg\,min}_{\begin{subarray}{c}\mathcal{S}\subset\mathcal{N}\\ |\mathcal{S}|=m\end{subarray}}\text{J}(\{o_{i}\}_{i\in\mathcal{S}};\theta)=\sum_{i\in\mathcal{S}}\left(f_{\theta}(o_{-i})-o_{i}\right)^{2}$ (7)
4.5 Free misalignment attacks
Our goal is to exploit alignment without access to observations. To illustrate our method, imagine a two-dimensional game in which two agents must jointly catch an object. Each agent observes the $(x,y)$ coordinates of the target. An adversary can manipulate their observations by pushing them toward completely different directions to prevent them from jointly catching the target. A suitable strategy is to push the agents in perpendicular directions [^1] We did not choose opposite directions as we cannot generalize the idea with more than 2 agents. We cannot have a set of three or more vectors where every pair points into opposite directions.. This idea generalizes to high-dimensional observations using vector orthogonality.
To design a misalignment attack, we will generate a matrix $\boldsymbol{\delta}=\begin{bmatrix}\delta_{1}^{\top}&\cdots&\delta_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{n\times d}$ that must satisfy two conditions:
  1. 1.
    Condition 1: The rows are orthogonal :
  2. 2.
    Condition 2: Satisfy the attack budget constraint
    $\|\delta_{i}\|_{\infty}\leq\epsilon\quad\forall i\in\mathcal{N}$ (9)
Generating such perturbations is nontrivial. Greedy construction can satisfy the budget constraint but often fails to ensure orthogonality, especially in high dimensions. Sampling and orthogonalizing a random matrix guarantees orthogonality but not the budget constraint (low magnitudes).
To generate perturbations that satisfy both constraints, we use partial Hadamard matrices $\tilde{\mathbf{H}}$ which are obtained by selecting a subset of rows from a full Hadamard matrix $\mathbf{H}$. Hadamard matrices have orthogonal rows, whose entries are either +1+1 or 1-1. Thus, it is clear that the following matrix satisfies both conditions (1) and (2):
$\boldsymbol{\delta}=\epsilon\times\tilde{\mathbf{H}}$ (10)
Hadamard matrices exist only for specific dimensions: when $d$ is a multiple of 4. To bypass this, we generate a full Hadamard matrix of size $\tilde{d}$, where $\tilde{d}$ is the largest power of two such that $\tilde{d}\leq d$:
$\tilde{d}=2^{\left\lfloor\log_{2}d\right\rfloor}$ (11)
We then pad the remaining columns with zeros. This padding affects neither orthogonality nor the budget constraint. We use Sylvester’s construction to generate Hadamard matrices.
5 Experiments
The goal of this section is to thoroughly investigate the performance of our attacks. To this end, we deliberately select multiple environments from three different MARL benchmarks, each chosen with a specific purpose to evaluate distinct aspects of our approach.
  • Level-Based Foraging (LBF) Papoudakis et al. (2021): is a grid-world environment where agents collect scattered food. We select fully observable, highly cooperative, and partially observable tasks. In fully observable tasks, agents observe the entire grid. In the highly cooperative setting (“-coop” in Figure 5), food is collected only if all agents surround it simultaneously, leading to overlapping observations that facilitate training $f_{\theta}$. Partially observable tasks are more challenging: agents see only a limited area around them (2-square or 1-square radius, with “-2s” or “-1s” flags). We use 10 LBF environments, with observation dimensions ranging from 15 to 18. The maximum number of agents is 4.
  • Multi-Robot Warehouse (RWARE) Papoudakis et al. (2021): is a partially observable multi-agent environment where robots are tasked with collecting requested shelves inside a warehouse. We test with three different levels of partial observability, where agents observe a $7\times 7$, $5\times 5$, or $3\times 3$ grid around them (“sr-3”, “sr-2” and “sr-1” flags in Figure 4). In addition to partial observability, this benchmark allows experiments with higher-dimensional observations, ranging from 71 to 351. The number of agents is 4 in all the tasks.
  • StarCraft Multi-Agent Challenge (SMAC) Ellis et al. (2024): SMAC is the most used benchmark in c-MARL and provides partially observable environments with a large number of agents, allowing us to evaluate how our algorithms scale. Since SMAC tasks are combat scenarios, agents can be removed during an episode due to death, leading to situations where the adversary loses access to some agents. This setting allows us to evaluate the robustness of our attacks under partial and dynamic access, where not all agents can be attacked throughout the episode. Such conditions reflect many realistic scenarios in which adversarial capabilities may be limited or intermittent. We selected six games with observation dimensions ranging from 82 to 285, and with 5, 10, or 27 agents.
Adversarial baselines: We use the following baselines: (1) White-box attack. (2) Random attacks. Most previous works considered only uniform and normal distributions. In addition, we include a non-symmetric distribution (exponential $\mathrm{Exp}(\lambda)$) and temporally correlated noise generated using the Ornstein–Uhlenbeck (OU) process. We report the performance of the best random attack (i.e, the random noise with the lowest return). Although comparing our limited-access attacks to a white-box attack is not entirely fair, we include it to estimate an upper bound and to select meaningful values for the attack budget as it is pointless to choose $\epsilon$ where the random attacks outperform the white‑box attack.
As we consider test-time attacks, we train the agents in each environment using QMIX  Rashid et al. (2018) and MAPPO Yu et al. (2022). We test different architectures to train $f_{\theta}$: feedforward, recurrent, and the encoder-only transformer. Unless otherwise stated, the results in the subsequent paragraphs are of agents trained with MAPPO, and the adversary collects a dataset of observations from 5,000 environment transitions and uses RNNS to train $f_{\theta}$. We fix PGD iterations at $K=10$ and the step size of $\alpha=\frac{\epsilon}{K}$. Full details of the used hyperparameters and additional results can be found in the appendix.
Metrics: We evaluate our attacks using episodic return. To ensure robust results, we report the interquartile mean (IQM) Agarwal et al. (2021) of returns over 50 independent episodes, along with 95% confidence intervals. IQM is robust to outliers and less biased than the median. Our code is open-source. [^2] Github: https://github.com/AmineAndam04/black-box-marl.git
5.1 Numerical results
Figures 34, and 5 present the IQM returns (y-axis) for SMAC, RWARE, and LBF benchmarks, respectively, across selected $\epsilon$ values (x-axis). The values at $\epsilon=0$ represent the benign performance of the agents during deployment.
- LBF (Figure 5): We observe that both Align and Hadamard attacks are effective in both fully and partially observable tasks (first two vs last three columns). Comparing each task with its highly cooperative variant (1st vs 2nd, and 3rd vs 4th columns), attacks are substantially more effective in highly cooperative settings, leading to much lower returns. Notably, in most partially observable scenarios, the Hadamard attack, despite being a free attack, matches or even outperforms the Align attack. The same pattern is observed on the RWARE benchmark.
- RWARE (Figure 4): Overall, both attacks are effective across tasks. Under the strongest partial observability (sr = 1, 1st and 4th columns), the Align attack remains effective but is outperformed by Hadamard. As partial observability decreases (e.g, from the 1st to the 3rd column), the performance gap narrows, as learning a good network $f_{\theta}$ becomes easier. Results on RWARE further indicate that attack effectiveness is not degraded by high dimensional observations.
- SMAC (Figure 3): The Align attack exhibits consistent performance across all games. In contrast, the Hadamard attack struggles to achieve significant gains over random noise in several scenarios (1st, 3rd, and 4th columns). Beyond partial observability and high dimensionality, SMAC further demonstrates that the Align attack remains effective in scenarios with large numbers of agents and without persistent access to the agents.
Next, we evaluate targeted attacks. For a fixed number of agents $m$, the Align attack uses the targeted procedure from Section 4.4, while Hadamard selects agents randomly. Results are shown in Figures 67, and 8, with the number of attacked agents indicated at the right of each subplot. Two observations stand out: first, Align maintains strong performance when 50% or more agents are targeted; second, Align is less sensitive than Hadamard to targeting fewer agents. To illustrate this second point, in Figures 7 and 8 we specifically show tasks where Hadamard outperforms Align when all agents are attacked, but as fewer agents are targeted, Align becomes more effective.
The last remark also indicates that the Align attack effectively identifies vulnerable agents. To further test this, we isolate the targeting mechanism: we select agents using the Align network (Equation 7) but inject Hadamard perturbations instead of Align perturbations.
In Tables 2 and 3, we report the performance gap between Hadamard attack obtained with and without selection mechanisms. For example, if the untargeted Hadamard attack causes a -15% drop relative to the benign return and the table reports a -10% drop, this means that the targeted Hadamard attack yields a total drop of -25%. Across all benchmarks, the targeted Hadamard attack consistently outperforms the untargeted variant, with a performance gap of more than two digits in many environments, with a maximum additional drop of -57%. On average across all the tasks, the additional drop is -11.5% on LBF, -6.18% on RWARE, and -6.28% on SMAC. These results demonstrate the effectiveness of the alignment network $f_{\theta}$ for target selection.
Figure 6: Performance of targeted attacks on SMAC.
Figure 7: Performance of targeted attacks on RWARE.
Figure 8: Performance of targeted attacks on LBF.
Task $m$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.25$
1s-10x10 2 -23.57 -4.99 -6.6
3 2.32 -6.28 -7.48
1s-8x8 1 -6.01 -0.96 -19.24
2 -7.42 -0.79 -19.34
2s-8x8-coop 1 0 -3.84 -3.84
2 -32.69 -48.07 -57.69
Table 2: Targeted Hadamard attack on LBF
Task $m$ $\epsilon=0.05$ $\epsilon=0.09$ $\epsilon=0.15$
tiny-4ag-sr-1 1 0.22 -0.37 -4.50
3 -1.35 -2.1 -16.80
small-4ag-sr-1 1 -9.80 -6.71 -2.35
3 -2.17 -14.51 -19.23
10m_vs_11m 3 -8.3 -10.17 -3.71
7 -14.35 0.40 1.98
MMM2 3 3.26 0.73 -6.32
7 -2.76 1.22 -7.10
Table 3: Targeted Hadamard attack on RWARE and SMAC
5.2 Ablations
Figure 9: Collected data size effect across MARL benchmarks
To normalize returns across tasks with different magnitudes, the y-axis in the following figures reports the percentage drop in return caused by our attack relative to the benign reward.
Data collection: Figure 9 evaluates the Align attack with training data ranging from 1,000 to 100,000 steps. Surprisingly, a strong attack performance is achieved with few samples, with no observable gains from larger datasets. This reduces data collection requirements, which improves the stealth of the attack and lowers storage and training costs for the adversary.
Network architecture: Figure 10 compares the Align attack across different neural architectures. Across all benchmarks, RNNs generally achieve the strongest performance, with a particularly large margin on RWARE. MLPs and Transformers perform comparably, but MLPs provide a better cost-performance trade-off due to faster training with only three hidden layers. RNNs share this advantage, as we employ shallow variants with one or two recurrent layers. We also tested deeper and more complex architectures, but they did not yield further performance gains.
Figure 10: Neural architecture effect across MARL benchmarks
PGD and FGSM: We investigate the impact of the number of PGD steps $K$. Figure 11 compares performance for $K\in\{1,5,10\}$. A key observation is that a single step is often sufficient to achieve strong performance, which is important for real-time attacks, as it allows perturbations to be generated quickly with minimal computational cost.
Figure 11: PGD’s iteration effect across MARL benchmarks.
5.3 Discussion
Most adversarial MARL work, including ours, evaluates attack effectiveness primarily using episodic return. We believe episode length is another important metric that is often overlooked, particularly in real-world scenarios with constrained adversaries. In such settings, fully incapacitating the c-MARL system is not always realistic; a more plausible outcome is delaying task completion by increasing episode length, even if tasks are eventually completed. Longer episodes can substantially increase operational costs, including energy and computation, making this a meaningful measure of adversarial impact. We illustrate this using the Align attack on LBF. Table 4 shows percentage increases in episode length, which often exceed double digits and reach up to 226% in the worst case.
Tasks $\epsilon=$ $0.1$ $0.15$ $0.2$ $0.25$ $0.5$
2s-8x8-coop 9.59 43.55 138.52 152.49 170.86
2s-10x10-coop 34.44 23.78 65.78 68.52 79.04
8x8-coop 10.59 29.67 103.86 188.56 226.80
8x8 5.94 21.06 51.93 75.93 193.21
Table 4: Episode length increase (%) on LBF tasks
Limitations: Our proposed algorithm faces the same core challenges as any MARL training procedure, such as partial observability, which we discuss thoroughly in the paper. In addition, issues like handling heterogeneous agents and multi-modal observations may further affect the effectiveness of our attack. Moreover, we studied realistic attacks from the perspective of the adversary’s access and capabilities. However, other perspectives can also be considered, including system-level constraints, attack deployment, and detectability in real-world settings. We plan to address these aspects in future work.
6 Conclusion
In this paper, we present two novel adversarial attacks specifically designed for collaborative MARL during deployment. We consider a black-box scenario under highly constrained conditions: First, a scenario where the adversary has access solely to the observations of deployed agents. Second, a more constrained scenario where the adversary has no access at all. Our approach exploits the reliance of collaborative agents on aligned perceptions for effective cooperation and adds small perturbations to the agents’ observations specifically designed to induce misaligned perceptions. Our empirical investigation showed the effectiveness of our attacks. Notably, our method requires minimal data to generate effective perturbations; we found that as few as 1,000 steps are sufficient to create impactful damage. Moreover, the adversary can choose to execute the attack in fewer steps while still achieving satisfactory results.
Appendix A supplementary material
A.1 Pseudo-codes
In Algorithm 1 and Algorithm 2 we provide the detailed pseudo-codes of align attack and PGD attack respectively.
@@ A.1 Pseudo-codes
A.2 Hadamard matrices
In our implementations, we generate Hadamard matrices using Sylvester’s construction. It is important to note that Hadamard matrices exist only for specific dimensions: when $d$ is a multiple of 4. To bypass the requirement that $d$ must be a multiple of 4, we generate a full Hadamard matrix of size $\tilde{d}$, where $\tilde{d}$ is the largest power of two such that $\tilde{d}\leq d$:
@@ A.2 Hadamard matrices
Figure 12: Padded partial Hadamard matrix for $n=3,d=10$.
@@
$\lambda\left\|\boldsymbol{\delta}^{\top}\boldsymbol{\delta}-\boldsymbol{I}\right\|_{2}^{2}.$ (13)
We implemented these two approaches, but did not notice any substantial improvements. However, using the targeted version of the Align attack to select a subset of agents and injecting Hadamard perturbations proved very effective in our experiments.
A.3 Experimental details
We provide full details on the used environmnets in table 5, 7, 6.
Environment FO PO HC $\text{dim}(o_{i})$ N
8x8-3p-2f 15 3
8x8-3p-2f-coop 15 3
1s-8x8-3p-2f 15 3
2s-8x8-3p-2f 15 3
2s-8x8-3p-2f-coop 15 3
10x10-4p-2f 18 4
10x10-4p-2f-coop 18 4
1s-10x10-4p-2f 18 4
2s-10x10-4p-2f 18 4
2s-10x10-4p-2f-coop 18 4
Table 5: Selected LBF environments: we select (FO) fully observable, (PO) partial observable, and (HC) highly cooperative tasks.
Environment PO SR $\text{dim}(o_{i})$ N
tiny-4ag-sr-1 $3\times 3$ 71 4
tiny-4ag-sr-2 $5\times 5$ 183 4
tiny-4ag-sr-3 $7\times 7$ 351 4
small-4ag-sr-1 $3\times 3$ 71 4
small-4ag-sr-2 $5\times 5$ 183 4
small-4ag-sr-3 $7\times 7$ 351 4
Table 6: Selected RWARE environments we select sensory range (SR) of $3\times 3$, $5\times 5$ and $7\times 7$.
Environment PO $\text{dim}(o_{i})$ N
27m_vs_30m 285 27
10m_vs_11m 105 10
MMM2 176 10
protoss_5_vs_5 92 5
terran_5_vs_5 82 5
zerg_5_vs_5 82 5
Table 7: Selected SMAC environments: all environments are partially observable (PO)
We trained QMIX and MAPPO agents using the hyperparameters specified in tables 8, 9, 10 , and LABEL:tab:mappo_rware_hyper. We use the Epymarl library Papoudakis et al. [2021] for training. We only report hyperparameters that are different from the default values found in Epymarl configurations. For the RWARE benchmark, we only train MAPPO agents, as QMIX performs poorly at this task Papoudakis et al. [2021]. Once the training is complete, we freeze the policy weights.
We test with three different neural architectures:
Table 8: QMIX Hyperparameters Used for SMAC and LBF benchmark
@@ A.3 Experimental details
Table 9: MAPPO Hyperparameters for SMAC benchmark.
@@
Table 10: MAPPO Hyperparameters for LBF benchmark.
@@
Table 11: MAPPO Hyperparameters for RWARE.
@@
tiny-4ag-sr-1 0.0005 False False 10 25M
tiny-4ag-sr-2 0.0008 False False 10 25M
tiny-4ag-sr-3 0.0008 False False 10 25M
small-4ag-sr-1 0.0005 False False 10 25M
small-4ag-sr-2 0.0008 False True 10 25M
small-4ag-sr-3 0.0008 False False 20 40M
A.4 Baselines:
We use the following baselines: